# Recognising perfect fits

### v1

@@ 3.2   Finding a free homotopy
Let $\mathcal{B}^{*}$ denote the cell decomposition of $M$ that is dual to $\mathcal{B}$. Let $G$ be a presentation of $\pi_{1}(M)$ computed from the cell structure of $\mathcal{B}^{*}$. An element of ¯$\overline{\mathcal{I}}_{n}$ gives a word in $G$. We use $\texttt{AreConjugate}(G,(w,v))$ to denote the solution to the conjugacy problem ([sela, Theorem 4.2], Préaux [preaux, Main Theorem]) applied to the presentation $G$ words $(w,v)$. We apply this solution to scan for pairs of itineraries in ¯$\overline{\mathcal{I}}_{n}$ whose encoded orbits are the corners of a common lozenge in $\mathcal{O}$. This gives $\mathtt{FindFit}(n,\mathcal{B})$, the following simple algorithm.
Algorithm 3.1 $\mathtt{FindFit}(\mathcal{B},n)$
@@
We use Fenley’s characterisation, 2.9. If $\varphi$ has perfect fits, there is a pair of (possibly the same orbit) that arises as the corners of a common lozenge in $\mathcal{O}$. The precise multiples that can occur (2.10) are precisely those encoded by the itineraries by 3.2. A pair or single orbit arises in ¯$\overline{\mathcal{I}}_{n}(\mathcal{B})$ for some $n$ by Lemma 3.4, at which point $\mathtt{FindFit}(n,\mathcal{B})$ returns $\mathtt{True}$. Since the itineraries in ¯$\overline{\mathcal{I}}_{n}(\mathcal{B})$ are unique, the converse follows similarly. ∎
@@
The periodic orbits of $\varphi(s)$ are precisely the images of the undrilled periodic orbits of $\varphi$ along with the cores of the filled solid tori. We then scan these orbits for conjugacies to see if $\varphi(s)$ has perfect fits in the classical sense, which is equivalent to $\varphi$ having no genuine perfect fits (2.12). ∎
@@ 4   Finding no perfect fits
Let $(\varphi,\{s_{i}\})$ be a (pseudo-)Anosov flow with marked orbits (2.4). Let $\mathcal{B}$ be a box decomposition of $(\varphi,\{s_{i}\})$. In this section, we give a routine to verify that a pseudo-Anosov $\varphi$ has no genuine perfect fits, Algorithm 3.1.
@@ 4.1   The Agol-Guéritaud construction
Let $r\subset\mathcal{P}$ be a rectangle with corresponding homeomorphism $f_{r}:r\to(0,1)^{2}$. Suppose that the boundary of $r$ is a union of cusp leaves. If:
@@
  • Each face rectangle contains three edge subrectangles
@@
  • Angles: If $e_{i}$ spans $r$ in either direction, assign $e_{i}$ a dihedral angle of $\pi$. Otherwise, assign $e_{i}$ a dihedral angle of 00.
@@
Now we assign face pairings between the tetrahedra $\{t(r)\}_{r\in\mathcal{R}}$. Let $f,f^{\prime}$ be a pair of face rectangles that are subrectangles of rectangles in $\mathcal{R}$. If $f$ and $f^{\prime}$ are $\pi_{1}(M^{\circ})$-translates, assign a pairing between the corresponding faces of $F$ and $F^{\prime}$. This pairing corresponds to extending $f$ into (some translate of) $r$ one way and $r^{\prime}$ the other in $\mathcal{P}$, see the right of Figure 8.
@@
4.2 is stated so that it can be applied to a finite collection $\mathcal{R}$ regardless of whether $\varphi$ has perfect fits.
@@ 4.2   Characterising no perfect fits
The flow $\varphi$ has no genuine perfect fits if and only if $\mathcal{P}$ admits a complete collection of tetrahedron rectangles.
@@
4.4 is essentially known in the literature, though in different forms. Consider first the forward direction. If $\varphi$ has no perfect fits, then $\mathcal{P}$ is loom in the sense of [ss-loom, Definition 2.11]. Then we apply the original Agol-Guéritaud construction following [ss-loom, §5] to obtain a locally finite veering triangulation. Since tetrahedron rectangles are finite up to the $\pi_{1}(M)$-action by [landry-minsky-taylor, Lemma 4.5], there is a finite complete collection $\mathcal{R}$. The fact that $\mathcal{P}$ is a loom space for a flow without perfect fits is (implicitly) asserted in [landry-minsky-taylor, §4], and a proof will appear in future work of Schleimer-Segerman’s program (promised in [ss-loom, Example 2.19]).
@@
¯=$\overline{\mathcal{R}}=\{g\cdot R:g\in\pi_{1}(M^{\circ})\}$
@@
Let $l\subset\mathcal{P}$ be a half of (that is, a ray contained in) a cusp leaf that meets $r\in\overline{\mathcal{R}}$. We cover $l$ with a layered sequence of tetrahedron rectangles $(r_{n})$, as follows. The finite end of $l$ is contained in some $r_{0}\in\overline{\mathcal{R}}$. There are two vertically spanning face subrectangles of $r_{0}$. At least one of these face rectangles $f_{0}\subset r_{0}$ contains $l\cap r_{0}$, with all three possible configurations shown in Figure 10.
@@
Now we choose $f_{0}$ as an unstable-spanning face rectangle containing $l\cap r_{0}$. If there is a choice, prioritise the face whose other (not $r_{0}$) tetrahedron rectangle meets a greater subset of $l$ than $r_{0}$ does. We choose a new rectangle $r_{i}$ inductively, giving us a layered sequence of tetrahedron rectangles $(r_{i})$. Note that the rectangles $r_{i}$ are strictly decreasing in the unstable direction and increasing in the stable direction. ∎
@@
For $l$ not a cusp leaf, there is a layered sequence $(r_{n})$ that covers $l$.
@@
Breaking symmetry, take $l$ to be the south half of a stable leaf. Since the sequence $(r_{n})$ decreases in the unstable direction, it suffices to show that there is a subsequence of $(r_{n})$ that is obtained by extending only over the south edges of the preceding rectangle. If this were not the case, the tail of $(r_{n})$ is obtained by only extending northward. This occurs only when $l$ meets each rectangle of the tail as in the rightmost of Figure 10. This implies that the rectangles must share a common edge subrectangle $e\subset\mathcal{P}$ in their south-east corners. The layered sequence will not terminate after finitely many rectangles, since $l$ is not a cusp leaf. This contradicts Lemma 4.5. ∎
@@
Let $q\in\mathcal{P}$ be any point. Then for any $p$ that is covered by $\overline{\mathcal{R}}$, there is a path from any $p\in r_{i}$ to $q$ given by a finite union of stable and unstable segments $\{I_{k}\}$. Lemma 4.7 implies that each $I_{k}$ can be covered by finitely many rectangles in $\overline{\mathcal{R}}$. So $q$ is contained in some rectangle of $\overline{\mathcal{R}}$. ∎
@@
Suppose that $\mathcal{P}$ admits a complete collection of tetrahedron rectangles $\mathcal{R}$. Then $\varphi$ has no genuine perfect fits.
@@
Suppose for contradiction that $\varphi$ has perfect fits. We again appeal to Lemma 4.5. Since $\varphi$ has perfect fits, there is a genuine perfect fit rectangle $Q\subset\mathcal{P}$ (2.8). Breaking symmetry, assume that the perfect fit occurs at the north-east corner of $Q$. Let $p\in Q$ be the south-east corner of $Q$. Since $\overline{\mathcal{R}}$ covers $\mathcal{P}$ (Lemma 4.8), there exists a rectangle $r_{0}\in\overline{\mathcal{R}}$ with $p\in r_{0}$.
Let $l\in\mathcal{P}^{s}$ be the stable leaf through $p$; this leaf forms the perfect fit in $Q$. Using 4.6, we build a layered sequence of rectangles $(r_{n}(l))$ that covers a half of $l$ forming the perfect fit, using $r_{0}(l)=r_{0}$. This sequence does not terminate, since this would give a tetrahedron rectangle with an ideal point that is not a cusp in its boundary. Let $k\in\mathcal{P}^{u}$ be the unstable leaf through $p$, on the south side of $Q$. The leaves $l$ and $k$ cut $r_{n}$ into quarters, and we refer to these quarters using the intercardinal directions.
Since the sequence $r_{i}$ decreases in the unstable direction, there is $N$ such that the western edge of $r_{N}$ meets the interior of $Q$. Since $Q$ is a perfect fit rectangle, there can be no cusps where $r_{N}$ meets $Q$. So, there are no cusps on the north-west side of $r_{N}$, as in the left of Figure 12.
Since $(r_{i})$ covers $l$, there is $N$ such that $r_{N+1}$ is obtained by extending over the north side of $r_{N}$. If there are no cusps in the south-east corner of $r_{N}$, choose $s_{0}=r_{N}$. If there are cusps in the south-east corner of $r_{N}$, we can perform one or two face extensions to $r_{N}$ (Figure 12, left) to obtain a rectangle $s_{0}\in\overline{\mathcal{R}}$ such that $s_{0}$ meets $l$ and $s_{0}$ does not have cusps in its south-east quarter. Now we build a new layered sequence starting from $s_{0}$ that covers the same end of $l$ as $r_{n}$. With this new sequence $(s_{n})$, we must always extend over the north side (see Figure 12, right). This gives a layered sequence that meets a fixed common south-west edge rectangle, which contradicts Lemma 4.5. ∎
@@
Question 4.11.
The argument in the proof of Lemma 4.8 describes how face pairings between tetrahedron rectangles must behave when there are perfect fit rectangles. However, skeletal rectangles can still exist when $\varphi$ has perfect fits. What is (if any) their significance? ∎
The main routine of the section, $\mathtt{FindVeering}$, will apply 4.2 and check the criterion 4.4. The remaining work is to build, from $\mathcal{B}$, the required skeletal rectangles. This comprises the remainder of the section.
@@ 4.3   Boxes in the universal cover
We introduce some nomenclature for working with $\mathcal{B}$. Declare two cusp rectangles in $\mathcal{P}$ as equivalent if their projections to $\mathcal{O}$ share a corner in $\{s_{i}\}$. A cusp of $\mathcal{P}$ is an equivalence class of cusp rectangles ([ss-loom, Definition 3.3]). Cusps in $\mathcal{P}$ lift naturally to cusps of the lifted flow space $N$. We also have stable/unstable cusp leaves of $N^{s}(\alpha)$/$N^{u}(\alpha)$ of $N$ as the lifts of cusp leaves in $\mathcal{P}$. In the other direction call, we call the quotient of some subset $X\subset N$ in $\mathcal{P}$ the shadow of $X$.
@@
Since $N$ is simply connected, any two boxes $A,B\in\mathcal{B}$ that are glued meet along a single two-cell. Call a gluing between $A$ and $B$ vertical if it occurs between the tops and bottoms of boxes, otherwise horizontal.
@@
(a) $A$ lies above $B$ lies above $C$
(b) $A$ is adjacent to $B$ and $C$ while $C$ lies above $B$
@@
Let $\alpha$ be a cusp of $N$. Let $L\in N^{s}(\alpha)$ be a stable (unstable) cusp leaf that emanates from $\alpha$ and into a box $B\in\mathcal{B}$. Suppose that $p\in L^{\alpha}$. Then there exists a box $A\in\mathcal{B}$ such that:
@@
  • the leaf $L$ emanates into $A$
@@
Breaking symmetry, take $L\in N^{s}$ a stable leaf. Let $\mathcal{A}$ be the collection of all boxes $A$ lying above $B$ such that $L$ emanates into $A$. As one walks upward through $\mathcal{A}$, the boxes must eventually stretch in the stable direction because of the Markov property (see 2.3). The wall containing $\alpha$ is a persistent unstable wall (3.3), so the boxes must extend over $L$. See Figure 14.
@@ 4.4   Finding rectangles
A cusp of $\mathcal{A}$ corresponds to some underlying cusp $\alpha$ of $N$, and we often use $\alpha$ to denote the corresponding cusp of $\mathcal{A}$. However, distinct cusps of $\mathcal{A}$ may correspond to the same cusp of $N$. Verifying otherwise runs into the same difficulty as in 4.14. However, it is trivial to see that two cusps are translates; it is when the marked orbit is the same. ∎
Define a partial order on $\mathcal{A}$ by declaring $B\leq A$ when there is an ascending vertical chain of boxes from $B$ to $A$ contained in $\mathcal{A}$ (this partial order is weaker than that inherited from $\mathcal{B}$). For a collection of boxes $\{B_{i}\}$ in $\mathcal{A}$, we use $\max_{i}(B_{i})$ to denote the set of maximal boxes in $\{B_{i}\}$ with respect to this partial order on $\mathcal{A}$. Given a cusp $(\alpha,\{B_{i}\})$, a stable leaf $L\in\mathcal{A}^{s}(\alpha)$ emanates into one or two boxes in $\max_{i}(B_{i})$. The Markov property ensures that $L$ emanates into the boxes of $\{B_{i}\}$ below $\max_{i}(B_{i})$. We use a single box to determine each cusp leaf by enforcing a tie-breaker.
@@
Fix a cusp $\alpha=(s,\{B_{i}\})\in\mathrm{Cusps}(\mathcal{A})$. Let $L\in N^{s}(\alpha)$. There is a unique box $B_{L}\in\mathrm{max}_{i}(\{B_{i}\})$ such that $L$ emanates into the interior or into the west wall of $B_{L}$. We call the pair $(B_{L},L)$ a stable cusp leaf of $\alpha$. We call $B_{L}$ the base box of the cusp leaf. Similarly, we define unstable cusp leaves, now using minimal boxes, and with leaves emanating into the interior or into the interior or the south wall of the boxes. We denote the stable and unstable cusp leaves of $\alpha$ as $\mathcal{A}^{s}(\alpha)$ and $\mathcal{A}^{u}(\alpha)$. ∎
@@
Without reference to a specific cusp, we use $\mathcal{A}^{s}$ and $\mathcal{A}^{u}$ to be the sets of all stable and unstable cusp leaves, respectively. Like cusps, we have potential duplicates, that is, distinct pairs $(B_{L},L)$, $(B^{\prime}_{L},L)\in\mathcal{A}^{s}$ for the same underlying leaf $L$.
@@
Let $(B_{L},L)\in\mathcal{A}^{s}$ be a stable (unstable) cusp leaf. Let $\mathcal{W}(A,(B_{L},L))$ be the collection of all boxes in $\mathcal{A}$ that lie below $B_{L}$ (above $B_{L}$). We call $\mathcal{W}(A,(B_{L},L))$ the wall of boxes about $L\in N^{s}$ in $\mathcal{A}$. ∎
@@
In $N$, the wall of boxes $\mathcal{W}(A,(B_{L},L))$ is a combinatorial neighbourhood of a connected component of where $L$ meets $\mathcal{A}$, see Figure 16.
Notation 4.19.
Unless the base box is required for clarity, we use an abuse of notation and refer to a cusp leaf $(B_{L},L)$ of $\mathcal{A}$ simply as the underlying leaf $L$. We say that a property holds for a cusp leaf $L\in\mathcal{A}^{s}$ when it holds for at least one pair of the form $(B_{L},L)\in\mathcal{A}^{s}$, and the same for unstable leaves. ∎
@@
Let $(L,K)\in\mathcal{A}^{s}\times\mathcal{A}^{u}$ be a pair of cusp leaves such that $\mathcal{W}(\mathcal{A},L)\cap\mathcal{W}(\mathcal{A},K)\neq\varnothing$. We call $(L,K)$ a joint, and let $\mathrm{Joints}(\mathcal{A})$ be the set of joints. ∎
@@
If $(L,K)\in\mathrm{Joints}(\mathcal{A})$, the pair of underlying leaves $L\in N^{s}$ and $K\in N^{u}$ intersect in $N$.
@@
If $(L,K)\in\mathrm{Joints}(\mathcal{A})$, there is $B\in\mathcal{W}(\mathcal{A},L)\cap\mathcal{W}(\mathcal{A},K)$. By the Markov property, $L$ and $K$ both meet $B$. So, $L$ and $K$ intersect in $B$. ∎
@@
Remark 4.22.
Suppose that there are distinct cusp leaves $(B_{L},L),(B_{L}^{\prime},L)\in\mathcal{A}^{s}$ that share an underlying leaf in $L\in N^{s}$. Let $(B_{K},K)\in\mathcal{A}^{u}$. Only one of the walls of boxes about $(B_{L},L)$ or $(B_{L}^{\prime},L)$ can share a box with the wall about $(B_{K},K)$. Otherwise, there would be a vertical chain of boxes connecting $B_{L}$ and $B_{L}^{\prime}$, so the cusp leaves are not distinct. So, there are effectively no duplicate cusp leaves when working with $\mathrm{Joints}(\mathcal{A})$, which justifies 4.19. ∎
@@
The forward direction follows immediately from Lemma 4.21. So, suppose that $L$ and $K$ intersect. Then they do so in some box $B\in\mathcal{B}$. By compactness of $M$, there is $n$ such that $B\in\mathcal{B}(n)$. By Lemma 4.12, there is $B_{L}\in\mathcal{B}$ lying above $B$ and $B_{K}\in\mathcal{B}$ below $B$ such that $L$ and $K$ emanate into $B_{L}$ and $B_{K}$ respectively. This implies that for $n$ sufficiently large, we have $B,B_{L},B_{K}\in\mathcal{B}(n)$. In turn, $B\in\mathcal{W}(\mathcal{B}(n),L)\cap\mathcal{W}(\mathcal{B}(n),K)$ and $(L,K)\in\mathrm{Joints}(\mathcal{B}(n))$. ∎

### v2

@@ 3.2  Finding a free homotopy
Let $\mathcal{B}^{*}$ denote the cell decomposition of $M$ that is dual to $\mathcal{B}$. Let $G$ be a presentation of $\pi_{1}(M)$ computed from the cell structure of $\mathcal{B}^{*}$. An element of $\overline{\mathcal{I}}_{n}$ gives a word in $G$. We use $\texttt{AreConjugate}(G,(w,v))$ to denote the solution to the conjugacy problem ([sela, Theorem 4.2], Préaux [preaux, Main Theorem]) applied to the presentation $G$ and words $(w,v)$. We apply this solution to scan for pairs of itineraries in $\overline{\mathcal{I}}_{n}$ whose encoded orbits are the corners of a common lozenge in $\mathcal{O}$. This gives $\mathtt{FindFit}(n,\mathcal{B})$, the following simple algorithm.
Algorithm 3.6 $\mathtt{FindFit}(n,\mathcal{B})$
@@
We use Fenley’s characterisation, 2.10. If $\varphi$ has perfect fits, there is a pair of (possibly the same orbit) that arises as the corners of a common lozenge in $\mathcal{O}$. The precise multiples that can occur (2.11) are precisely those encoded by the itineraries by 3.2. A pair or single orbit arises in $\overline{\mathcal{I}}_{n}(\mathcal{B})$ for some $n$ by Lemma 3.5, at which point $\mathtt{FindFit}(n,\mathcal{B})$ returns $\mathtt{True}$. Since the itineraries in $\overline{\mathcal{I}}_{n}(\mathcal{B})$ are unique, the converse follows similarly. ∎
@@
The periodic orbits of $\varphi(s)$ are precisely the images of the undrilled periodic orbits of $\varphi$ along with the cores of the filled solid tori. We then scan these orbits for conjugacies to see if $\varphi(s)$ has perfect fits. This is equivalent to $(\varphi,\{s_{i}\})$ having no genuine perfect fits by 2.13. ∎
@@ 4 Finding no perfect fits
Let $(\varphi,\{s_{i}\})$ be a (pseudo-)Anosov flow with marked orbits (2.4). Let $\mathcal{B}$ be a box decomposition of $(\varphi,\{s_{i}\})$. In this section, we give a routine to verify that $(\varphi,\{s_{i}\})$ has no genuine perfect fits, Algorithm 3.6. The example to keep in mind is $\varphi$ a pseudo-Anosov flow and $\{s_{i}\}$ the set of singular orbits.
@@ 4.1  The Agol-Guéritaud construction
A cusp leaf of $\mathcal{P}$ is a lift of a punctured leaf in $\mathcal{O}^{\circ}$. Let $r\subset\mathcal{P}$ be a rectangle with corresponding homeomorphism $f_{r}:r\to(0,1)^{2}$. Suppose that the boundary of $r$ is a union of cusp leaves. If:
@@
  • Each face rectangle contains three edge subrectangles.
@@
  • Angles: If $e_{i}$ spans $r$ in either direction, assign $e_{i}$ a dihedral angle of $\pi$. Otherwise, assign $e_{i}$ a dihedral angle of 0.
@@
Now we assign face pairings between the tetrahedra $\{t(r)\}_{r\in\mathcal{R}}$. Let $f,f^{\prime}$ be a pair of face rectangles that are subrectangles of rectangles $r,\,r^{\prime}\in\mathcal{R}$. If $f$ and $f^{\prime}$ are $\pi_{1}(M^{\circ})$-translates, assign a pairing between the corresponding faces of $F$ and $F^{\prime}$. This pairing corresponds to extending $f$ into $r$ one way and into (some translate of) $r^{\prime}$ the other, see the right of Figure 8.
@@
We emphasise that 4.2 is stated so that it can be applied to a finite collection $\mathcal{R}$ regardless of whether $(\varphi,\{s_{i}\})$ has genuine perfect fits.
@@ 4.2  Characterising no perfect fits
The flow with marked orbits $(\varphi,\{s_{i}\})$ has no genuine perfect fits if and only if $\mathcal{P}$ admits a complete collection of tetrahedron rectangles.
@@
4.4 is essentially known in the literature, though in different forms. Consider first the forward direction. If $(\varphi,\{s_{i}\})$ has no perfect fits, then $\mathcal{P}$ is loom in the sense of [ss-loom, Definition 2.11]. Then we apply the original Agol-Guéritaud construction following [ss-loom, §5] to obtain a locally finite veering triangulation. Since tetrahedron rectangles are finite up to the $\pi_{1}(M)$-action by [landry-minsky-taylor, Lemma 4.5], there is a finite complete collection $\mathcal{R}$. The fact that $\mathcal{P}$ is a loom space for a flow without genuine perfect fits is (implicitly) asserted in [landry-minsky-taylor, §4], and a proof will appear in future work of Schleimer-Segerman’s program (promised in [ss-loom, Example 2.19]).
@@
$\overline{\mathcal{R}}=\{g\cdot R:g\in\pi_{1}(M^{\circ})\}$
@@
Let $l\subset\mathcal{P}$ be a half of (that is, a ray contained in) a leaf that meets $r\in\overline{\mathcal{R}}$. We cover $l$ with a layered sequence of tetrahedron rectangles $(r_{n})$, as follows.
Breaking symmetry, suppose that $l$ is half of a stable leaf extending southward. The finite end of $l$ is contained in some $r_{0}\in\overline{\mathcal{R}}$. There are two vertically spanning face subrectangles of $r_{0}$. At least one of these face rectangles $f_{0}\subset r_{0}$ contains $l\cap r_{0}$, with all three possible configurations shown in Figure 10.
@@
If there is a choice for $f_{0}$, then one of the choices can be extended over its south edge to a tetrahedron rectangle. Take this as $f_{0}$. With $f_{0}$ chosen, let $r_{1}$ be the other tetrahedron rectangle (not $r_{0}$) that $f_{0}$ spans. We again have a choice of unstable-spanning face rectangle $f_{1}$ of $r_{1}$, and proceed inductively, always preferring the face rectangle $f_{i}$ which extends southward when given a choice. This gives a layered sequence of tetrahedron rectangles $(r_{i})$. Note that the rectangles $r_{i}$ are strictly decreasing in the unstable direction and increasing in the stable direction.
If $l$ is the end of a cusp leaf which meets a cusp, then the sequence $(r_{i})$ can terminate with a rectangle that has this cusp on an edge. Otherwise, the sequence is infinite. ∎
@@
There is a layered sequence $(r_{n})$ that covers $l$.
@@
Breaking symmetry, take $l$ to be the south half of a stable leaf. Since the sequence $(r_{n})$ decreases in the unstable direction, it suffices to show that there is a subsequence of $(r_{n})$ that is obtained by extending only over the south edges of the preceding rectangle. If this were not the case, the tail of $(r_{n})$ is obtained by only extending northward. This occurs only when $l$ meets each rectangle of the tail as in the rightmost of Figure 10. This implies that the rectangles must share a common edge subrectangle $e\subset\mathcal{P}$ in their south-east corners.
If the layered sequence terminates after finitely many rectangles, then $l$ must be the side of cusp leaf meeting a cusp. In such a case, the rectangles then cover $l$.
So, suppose that $l$ does not end in a cusp. Then the sequence of rectangles $(r_{n})$ does not terminate. This contradicts Lemma 4.5. ∎
@@
By Lemma 4.7, $\overline{\mathcal{R}}$ covers every leaf of $\mathcal{P}$. Since every point lies on a leaf, $\overline{\mathcal{R}}$ covers $\mathcal{P}$. ∎
@@
Suppose that $\mathcal{P}$ admits a complete collection of tetrahedron rectangles $\mathcal{R}$. Then $(\varphi,\{s_{i}\})$ has no genuine perfect fits.
@@
Suppose for contradiction that $(\varphi,\{s_{i}\})$ has perfect fits. Then there is a genuine perfect fit rectangle $Q\subset\mathcal{P}$ (2.8). Breaking symmetry, assume that the perfect fit occurs at the north-east corner of $Q$. Let $p\in Q$ be the south-east corner of $Q$. Since $\overline{\mathcal{R}}$ covers $\mathcal{P}$ (Lemma 4.8), there exists a rectangle $r_{0}\in\overline{\mathcal{R}}$ with $p\in r_{0}$.
Let $l\in\mathcal{P}^{s}$ be the stable leaf through $p$; this leaf forms the perfect fit in $Q$. Using 4.6, we build a layered sequence of rectangles $(r_{n}(l))$ that covers a half of $l$ forming the perfect fit, using $r_{0}(l)=r_{0}$. This sequence does not terminate, since the half of $l$ we are covering does not contain a cusp. Let $k\in\mathcal{P}^{u}$ be the unstable leaf through $p$, on the south side of $Q$. The leaves $l$ and $k$ cut $r_{n}$ into quarters, and we refer to these quarters using the intercardinal directions.
Since the sequence $r_{i}$ decreases in the unstable direction, there is $N$ such that the western edge of $r_{N}$ meets the interior of $Q$. Since $Q$ is a perfect fit rectangle, there can be no cusps where $r_{N}$ meets $Q$. So, there are no cusps on the north-west quarter of $r_{N}$, as in the left of Figure 12.
Since $(r_{i})$ covers $l$, there is $N$ such that $r_{N+1}$ is obtained by extending over the north edge of $r_{N}$. If there are no cusps in the south-east quarter of $r_{N}$, choose $s_{0}=r_{N}$. If there are cusps in the south-east quarter of $r_{N}$, we can perform one or two face extensions to $r_{N}$ (Figure 12, left) to obtain a rectangle $s_{0}\in\overline{\mathcal{R}}$ such that $s_{0}$ meets $l$ and $s_{0}$ does not have cusps in its south-east quarter. Now we build a new layered sequence starting from $s_{0}$ that covers the same end of $l$ as $r_{n}$. With this new sequence $(s_{n})$, we must always extend over the north edge (see Figure 12, right). This gives a layered sequence that meets a fixed common south-west edge rectangle, which contradicts Lemma 4.5. ∎
@@
The main routine of the section, $\mathtt{FindVeering}$, will apply 4.2 and check the criterion $4.$ The remaining work is to build, from $\mathcal{B}$, the required skeletal rectangles. This comprises the remainder of the section.
@@ 4.3  Boxes in the universal cover
We introduce some nomenclature for working with $\mathcal{B}$. Declare two cusp rectangles in $\mathcal{P}$ as equivalent if their projections to $\mathcal{O}$ share a corner in $\{s_{i}\}$. A cusp of $\mathcal{P}$ is an equivalence class of cusp rectangles ([ss-loom, Definition 3.3]). Cusps in $\mathcal{P}$ lift naturally to cusps of the lifted flow space $N$. For $\alpha$ a cusp of $N$, we have the stable/unstable cusp leaves of $N$ as the lifts of the cusp leaves in $\mathcal{P}$. We use $N^{s}(\alpha)$ and $N^{u}(\alpha)$ to denote the set of all stable/unstable cusp leaves through $\alpha$. We use $N^{s}$ and $N^{u}$ to denote set of stable and unstable leaves of $N$, respectively.
In the other direction, we call the quotient of some subset $X\subset N$ in $\mathcal{P}$ the shadow of $X$.
@@
Call a gluing between $A$ and $B$ vertical if it occurs between the tops and bottoms of boxes, otherwise horizontal.
@@
(a) AA lies above $B$ lies above $C$
(b) AA is adjacent to $B$ and $C$ while $C$ lies above $B$
@@
Let $\alpha$ be a cusp of $N$. Let $L\in N^{s}(\alpha)$ be a stable (unstable) cusp leaf that emanates from $\alpha$ and into a box $B\in\mathcal{B}$. Suppose that $p\in L$. Then there exists a box $A\in\mathcal{B}$ such that:
@@
  • the leaf $L$ emanates into AA
@@
Let $\mathcal{A}$ be the collection of all boxes $A$ lying above $B$ such that $L$ emanates into $A$. As one walks upward through $\mathcal{A}$, the boxes must eventually stretch in the stable direction because of the Markov property (see 2.3). So, the boxes must extend over $L$. See Figure 14.
@@ 4.4  Finding rectangles
A cusp of $\mathcal{A}$ corresponds to some underlying cusp $\alpha$ of $N$, and we often use $\alpha$ to denote the corresponding cusp of $\mathcal{A}$. However, distinct cusps of $\mathcal{A}$ may correspond to the same cusp of $N$. Verifying otherwise runs into the same difficulty as in 4.13. However, it is trivial to check whether two corresponding cusps of $N$ are translates: it is when the marked orbit is the same. ∎
Define a partial order on $\mathcal{A}$ by declaring $B\leq A$ when there is an ascending vertical chain of boxes from $B$ to $A$ contained in $\mathcal{A}$ (this partial order is weaker than that inherited from $\mathcal{B}$). For a collection of boxes $\{B_{i}\}$ in $\mathcal{A}$, we use $\max_{i}(B_{i})$ to denote the set of maximal boxes in $\{B_{i}\}$ with respect to this partial order on $\mathcal{A}$. Given a cusp $(\alpha,\{B_{i}\})$, a stable leaf $L$ meeting this cusp emanates into one or two boxes in $\max_{i}(B_{i})$. For $B\in\max_{i}(B_{i})$, the Markov property ensures that $L$ emanates into each box in a vertical ascending chain up to $B$. We use a single box to determine each cusp leaf by enforcing a tie-breaker.
@@
Fix a cusp $\alpha=(s,\{B_{i}\})\in\mathrm{Cusps}(\mathcal{A})$. Let $L\in N^{s}(\alpha)$. There is at most one box $B_{L}\in\mathrm{max}_{i}(B_{i})$ such that $L$ emanates into the interior or into the west wall of $B_{L}$. If there is such a box, call the pair $(B_{L},L)$ a stable cusp leaf of $\alpha$. We call $B_{L}$ the base box of the cusp leaf. Similarly, we define unstable cusp leaves, now using minimal boxes, and with leaves emanating into the interior or into the south wall of the boxes. ∎
@@
We use $\mathcal{A}^{s}$ and $\mathcal{A}^{u}$ to denote the sets of all stable and unstable cusp leaves, respectively. Like cusps, we have potential duplicates, that is, distinct pairs $(B_{L},L)$, $(B^{\prime}_{L},L)\in\mathcal{A}^{s}$ for the same underlying leaf $L$.
@@
Let $(B_{L},L)\in\mathcal{A}^{s}$ be a stable (unstable) cusp leaf. Let $\mathcal{W}(\mathcal{A},(B_{L},L))$ be the collection of all boxes in $\mathcal{A}$ that lie below $B_{L}$ in $\mathcal{A}$ (above $B_{L}$). We call $\mathcal{W}(\mathcal{A},(B_{L},L))$ the wall of boxes about $(B_{L},L)$ in $\mathcal{A}$. ∎
@@
In $N$, the wall of boxes $\mathcal{W}(\mathcal{A},(B_{L},L))$ is a combinatorial neighbourhood of a connected component of where $L$ meets $\mathcal{A}$, see Figure 16.
@@
Let $(L,K)\in N^{s}\times N^{u}$ be a pair of cusp leaves. Call $(L,K)$ a joint if $\mathcal{W}(\mathcal{A},(B_{L},L))\cap\mathcal{W}(\mathcal{A},(B_{K},K))\neq\varnothing$ for some $(B_{L},L)$,  $(B_{K},K)$. We use $\mathrm{Joints}(\mathcal{A})$ to denote the set of joints. ∎
@@
If $(L,K)\in\mathrm{Joints}(\mathcal{A})$, the pair of leaves $L$ and $K$ intersect in $N$.
@@
If $(L,K)\in\mathrm{Joints}(\mathcal{A})$, there is $B\in\mathcal{W}(\mathcal{A},(B_{L},L))\cap\mathcal{W}(\mathcal{A},(B_{K},K))$. By the Markov property, $L$ and $K$ both meet $B$. So, $L$ and $K$ intersect in $B$. ∎
@@
The forward direction follows immediately from Lemma 4.19. So, suppose that $L$ and $K$ intersect. Then they do so in some box $B\in\mathcal{B}$. By compactness of $M$, there is $n$ such that $B\in\mathcal{B}(n)$. By Lemma 4.11, there is $B_{L}\in\mathcal{B}$ lying above $B$ and $B_{K}\in\mathcal{B}$ below $B$ such that $L$ and $K$ emanate into $B_{L}$ and $B_{K}$ respectively. This implies that for $n$ sufficiently large, we have $B,B_{L},B_{K}\in\mathcal{B}(n)$. In turn, $B\in\mathcal{W}(\mathcal{B}(n),(B_{L},L))\cap\mathcal{W}(\mathcal{B}(n),(B_{K},K))$ and $(L,K)\in\mathrm{Joints}(\mathcal{B}(n))$. ∎